# From Maximum Cut to Maximum Independent Set

### v1

@@ III.3.4 Special Benchmarks
Now we turn to deterministic instances. In a recent review article on MIS Marino-2024 , several popular benchmarks used in the literature are provided. Among those benchmarks, the one arising from coding theory Sloane-2000 is considered to be very challenging. We now use this benchmark to further assess the performance of those algorithms discussed previously. The detailed information of the instances in this benchmark is list in Table 2.
@@
From the results in Table 3 one can see that the performance of the three algorithms differ a lot. On average, CirCut performs better than MIN, especially for these three cases: 1dc.512, 1dc.1024 and 1dc.2048 (Here 1dc refers to Single-Deletion-Correcting Codes Sloane-2000 ; Sloane-2002 ). The performance of SA is even better than CirCut. Specifically, SA obtains the optimal values for all the instances whose independence number is known. And for the remaining four graphs, namely 1dc.2048, 1zc.1024, 1zc.2048 and 1zc.4096, the predictions from SA coincide with the lower bounds known. In essence, the best results from SA are exactly the same as those from a local search strategy long ago Butenko-2009 . This gives more confidence that these lower bounds could actually be optimal.
Furthermore, SA could provide additional information for the corresponding codes. Let us check the results for the 1dc graphs more carefully. It is known Sloane-2000 ; Sloane-2002 that, for $n\leq 10$ the Varshamov-Tenegolts code VT-1965 $\mbox{VT}_{0}(n)$ are optimal for these graphs. However, they are not the unique solutions. For example, one may replace the codes $(110100)_{2}$ and $(001011)_{2}$ in $\mbox{VT}_{0}(6)$ by $(111000)_{2}$ and $(000111)_{2}$, and obtain another MIS Sloane-2002 . Our solutions suggest that we can do the same for $\mbox{VT}_{0}(10)$, by replacing the corresponding decimal codewords $95$ and $928$ with $63$ and $960$. Moreover, we know $\mbox{VT}_{0}(11)$ gives the lower bound $172$ of the independence number of 1dc.2048 Sloane-2000 . Our calculation produces another independent set of size $172$, but it contains almost no common vertices with $\mbox{VT}_{0}(11)$.

### v2

@@ III.3.4 Special Benchmarks
Now we turn to deterministic instances. In a recent review article on MIS Marino-2024 , several popular benchmarks used in the literature are provided. Among those benchmarks, the one arising from coding theory Sloane-2000 is considered to be very challenging. We now use this benchmark to further assess the performance of those algorithms discussed previously. The detailed information of the instances in this benchmark is list in Table 2. Notice that the independence number of 1dc.2048 has recently been confirmed to be $172$ No-2019 ; Nakasho-2023 , and the upper bound for 1dc.4096 has been lowered to $320$ No-2019 .
@@
From the results in Table 3 one can see that the performance of the three algorithms differ a lot. On average, CirCut performs better than MIN, especially for these four cases: 1dc.512, 1dc.1024, 1dc.2048 and 1dc.4096 (Here 1dc refers to Single-Deletion-Correcting Codes Sloane-2000 ; Sloane-2002 ). The performance of SA is even better than CirCut. Specifically, SA obtains the optimal values for all the instances whose independence number is known. And for the remaining four graphs, namely 1dc.4096, 1zc.1024, 1zc.2048 and 1zc.4096, the predictions from SA coincide with the lower bounds known. In essence, the best results from SA are exactly the same as those from a local search strategy long ago Butenko-2009  (except for 1dc.4096 which is not calculated in Butenko-2009 ). This gives more confidence that these lower bounds could actually be optimal.
Furthermore, SA could provide additional information for the corresponding codes. Let us check the results for the 1dc graphs more carefully. It is known Sloane-2000 ; Sloane-2002 that, for $n\leq 10$ the Varshamov-Tenegolts code VT-1965 $\mbox{VT}_{0}(n)$ are optimal for these graphs. However, they are not the unique solutions. For example, one may replace the codes $(110100)_{2}$ and $(001011)_{2}$ in $\mbox{VT}_{0}(6)$ by $(111000)_{2}$ and $(000111)_{2}$, and obtain another MIS Sloane-2002 . Our solutions suggest that we can do the same for $\mbox{VT}_{0}(10)$, by replacing the corresponding decimal codewords $95$ and $928$ with $63$ and $960$. Moreover, we now know $\mbox{VT}_{0}(11)$ gives a MIS of 1dc.2048 Sloane-2000 ; No-2019 ; Nakasho-2023 . Our calculation produces another independent set of size $172$, and it contains almost no common vertices with $\mbox{VT}_{0}(11)$.
@@ IV Summary
Note added:
After finishing the paper, we learn that ref. Krpan-2024 has already studied the choice of the penalty parameter $\lambda$ and the post-processing procedure in great detail. The preferred choice $\lambda=1$ in Krpan-2024 coincides with ours. It would be interesting to further investigate the benefits of the post-processing procedure proposed in Krpan-2024 .